# Transformer-Based Rate Prediction for Multi-Band Cellular Handsets

### v1

@@ Abstract
Cellular wireless systems are witnessing the proliferation of frequency bands over a wide spectrum, particularly with the expansion of new bands in FR3. These bands must be supported in user equipment (UE) handsets with multiple antennas in a constrained form factor. Rapid variations in channel quality across the bands from motion and hand blockage, limited field-of-view of antennas, and hardware and power-constrained measurement sparsity pose significant challenges to reliable multi-band channel tracking. This paper formulates the problem of predicting achievable rates across multiple antenna arrays and bands with sparse historical measurements. We propose a transformer-based neural architecture that takes asynchronous rate histories as input and outputs per-array rate predictions. Evaluated on ray-traced simulations in a dense urban micro-cellular setting with FR1 and FR3 arrays, our method demonstrates superior performance over baseline predictors, enabling more informed band selection under realistic mobility and hardware constraints.
@@ I Introduction
Predictors enable proactive switching and lightweight channel inference: AE+LSTM sequence models improve Top-$C$ ranking and reduce misalignment for next-slot beam/BS prediction [16]; cross-band learning maps sub-6  GHz features to mmWave beam/blockage; vision/semantics aid LoS/NLoS forecasting and candidate selection [17, 18, 19]. Physics-informed Reinforcement Learning (RL) further embeds channel dynamics, demonstrating zero-shot and DT-in-the-loop gains with growing industrial traction [20, 21, 22].
@@ 3.5 GHz (100MHz)
For logging and a baseline, we use an exponential-weights forecaster (EXP3/Hedge) with $\varepsilon$-greedy exploration [28]. After collecting ray tracing data, we run an $\varepsilon$-greedy exponential-weights bandit ($\varepsilon=0.2$) to choose at each step a hopping observation subset of RX antennas based on historical scores. The chosen subset defines a mask: the UE keeps the corresponding ray-traced capacities and treats the rest as missing. We log the masked observations and the chosen subsets, and use the full capacities as labels to train the rate-only predictor. The generated dataset, including 1000 input instances, is shuffled and divided into 70%:15%:15% for training, validation, and testing. The whole project can be found in GitHub repository.

### v2

@@ Abstract
Cellular wireless systems are facing a proliferation of frequency bands over a wide spectrum, particularly with the expansion into FR3. These bands must be supported in user equipment (UE) handsets with multiple antennas in a constrained form factor. Rapid variations in channel quality across the bands from motion and hand blockage, limited field-of-view of antennas, and hardware and power-constrained measurement sparsity pose significant challenges to reliable multi-band channel tracking. This paper formulates the problem of predicting achievable rates across multiple antenna arrays and bands with sparse historical measurements. We propose a transformer-based neural architecture that takes asynchronous rate histories as input and outputs per-array rate predictions. Evaluated on ray-traced simulations in a dense urban micro-cellular setting with FR1 and FR3 arrays, our method demonstrates superior performance over baseline predictors, enabling more informed band selection under realistic mobility and hardware constraints.
@@ I Introduction
Predictors enable proactive switching and lightweight channel inference: AE+LSTM sequence models improve Top-$C$ ranking and reduce misalignment for next-slot beam/BS prediction [19]; cross-band learning maps sub-6  GHz features to mmWave beam/blockage; vision/semantics aid LoS/NLoS forecasting and candidate selection [3, 25, 26]. Physics-informed Reinforcement Learning (RL) further embeds channel dynamics, demonstrating zero-shot and DT-in-the-loop gains with growing industrial traction [27, 12, 13, 11].
@@ 3.5 GHz (100MHz)
For logging and a baseline, we use an exponential-weights forecaster (EXP3/Hedge) with $\varepsilon$-greedy exploration [4]. After collecting ray tracing data, we run an $\varepsilon$-greedy exponential-weights bandit ($\varepsilon=0.2$) to choose at each step a hopping observation subset of RX antennas based on historical scores. The chosen subset defines a mask: the UE keeps the corresponding ray-traced capacities and treats the rest as missing. We log the masked observations and the chosen subsets, and use the full capacities as labels to train the rate-only predictor. The generated dataset, including 1000 input instances, is shuffled and divided into 70%:15%:15% for training, validation, and testing. The whole project can be found in https://github.com/Ruibin2000/UE_fc_hopping_v2.git.